UT Mar 05 2022 - Split Night PETS/PEPSI - DEC



Night Info

  • Sunset 18:19
  • Nautical Twilight Ends 19:13
  • Astronomical Twilight Ends 19:41
  • Moon sets 20:39
  • Handover to PEPSI/PETS 21:10
  • PETS-7 21:40-02:02
  • Handover from PEPSI/PETS 02:03
  • Astronomical Twilight Starts 05:21
  • Nautical Twilight Stars 05:49
  • Sunrise 06:43

Staff

  • OSA: J. Williams
  • ISp: E.Solheid
  • Mtn Manager: Rick Hansen
  • ISA/Observer: B.Rothberg, R.T.Gatto
  • AO: G. Taylor

Night Plan

  • Install ao-supervisor update for testing (before sunset)
  • Starting Config LBCs
  • Twilight Flats if clear
  • TMS testing (both sides active) twilight-21:10
    • Halley's Comet 08 15 31.96 +02 27 40.0
  • LUCI-1 AO/ESM Testing 02:18-twlight
    • /home/lbto/Service_Observing/2022A_D_E/LUCI/AO_ESM/
      • AO OT Version 75
      • ESM OT Version 36
    • Bin 2 ESM w/new AO Supervisor
      • M67 - twilight-02hr
      • Dual BLR - 02hr-twilight
    • Close separation targets (0.1"-1")
      • INAF Target J0732+3533 - 2" separation 1.5mag difference - could test this script on-sky beforehand -its vsisble first part of night (could roll over to Sat night)
      • BD+33 2137 RA10-11hr Star 1 in Binary folder R=10.3 - separation 2.4" separation 0.5 mag flux difference (might be too far apart - but wiki log showed this was an issue - but seeing was unsteady)
      • RA12h-13hr Star 1 in Binary Folder BD+00 2888 R=10.7 - separation 0.225"
      • HIP 65069 - Ra13hr+17Dec R=6.9 0.12" separation - need 64x2048 sub-frame (use to also test sub-framing)
      • 2MASSJ1128+3136 R=13.5 - 1" separation - 1mag difference in flux
      • 2MASSJ13061537+2043444 R=8.5 1.6" separation (64x2048 sub-framing) - 1.2mag difference in flux
    • Bin 3 IQ
      • bright
      • faint
    • Check on Bin 2 Full AO IQ
    • Continue to test RTD Sub-frame
    • Verify no jumps in AO w/LUCI N30 - IT 8578/IT8579
      • Repeat Observing RA14hr-15hr Star 1 R=10.2 BD +00 3109 (saw jump on this before)
      • S255 (UVa target) - twilight-midnight
    • Finish on Cat's Eye Nebula (FeII and Pa-Beta if possible)

Poor Weather Backup

  • LUCI PhotStd - if clear and poor seeing (2" or greater - can convert to track mode)
    • /home/lbto/Service_Observing/2022A_D_E/PoorWeather/LUCI_Phot_Standards
  • Rothberg Mergers
    • OT: /home/lbto/Service_Observing/2022A_D_E/PoorWeather/Rothberg/LocalMergers/COBandhead_Mergers_V41.xml
      • NGC 4004 - 21hr-twilight
      • NGC 3597 - 23hr-02hr
      • UGC 9829 - 01:30hr-twilight
      • UGC 10607 - 02hr -twilight
      • UGC 10675 - 02hr-twilight

Night Log

Closed Dome

  • UT 00:01 - starting LBCs - will take test biases
  • UT 00:06 - bringing up LUCIs - noticed that LUCI-1 has unknown for Camera, will init, unknown on LUCI-2 init
    • Stoopid Hooman Error MCE/ROE wasn't powered on - I naviely assumed it was ON - Dave and I turned it on - its all good everything init ok'd
  • UT 00:35 - now taking LBC test biases
  • UT 00:36 - GEIRS error - I think its becasuse I turned on MCE/ROE after restarting service, so restarting GEIRS
    • yep - lots of persistence on LUCI-1 - this is because the power outage occurred and the J-band was in place.
  • UT 00:39 - LBC biases look OK
  • UT 00:46 - N30 Fieldstop aligned
  • UT 00:47 - D.Thompson restarting the LUCI software which includes PFC jump fix
  • UT 00:48 - winds at 30m/s - thick clouds coming in and out - probably not opening at sunset
  • UT 01:40 - O.Kuhn copied the last TMS reference, from 20220303, to /tmp (as telescope@robs)
  • UT 01:45 - winds still hitting limits, 29.5 m/s max, thick clouds to south
  • UT 02:25 - switched to test ao-supervisor, J.C. updated Bin 2 and Bin 3 tables
  • UT 02:41 - winds downtrending - still have 23m/s bumps, clouds have cleared out (for the most part) but the humidity is now 87%
  • UT 03:05 - now at 90% humidity, winds 25m/s
  • UT 03:28 - little while ago crossed the 95% threshold - winds 23m/s, thick clouds keep rolling in and out
    • ALLSKY camera looks a bit darker - contrast issue? IT 8586
    • moire artifacting seems to be increasing
  • UT 04:06 - reconfiguring for PEPSI-Bino 4min earlier
  • UT 04:17 - powering off LBCs

PEPSI/PETS

  • UT 05:24 - winds are under 20m/s but 99% humidity and the air temp and dewpoint are equivalent
    • J.Williams is checking if LBTO is in popsicle mode
    • rime ice forming
    • J. Williams is calling the night, he's calling it "Fred"
We are done....Handing over to D.Miller for AO closed dome testing.

-- BarryRothberg - 28 Feb 2022
Topic revision: r8 - 05 Mar 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback